OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the point (5, −3) and has a slope of −1.25. y = x +

OpenStudy (anonymous):

Please help me

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (cwrw238):

use the general form y-y1 = m(x-x1) where m is the slope and (x1,y1) is a point on the line

OpenStudy (anonymous):

Okay

OpenStudy (cwrw238):

so you plug in x = 5,y = -3 and m = -1.25 and simpify

OpenStudy (anonymous):

Im sorry im bad at math Do i need to divide ??

OpenStudy (anonymous):

Im stupid im sorry but i am

OpenStudy (cwrw238):

y - (-3) = -1.25(x - 5) i just replaced x ,y and m by -3,5 and -1.25

OpenStudy (cwrw238):

that gives y + 3 = -1.25x + 6.25 now subtract 3 from both sides of the '=' to get the equation in the form you require

OpenStudy (anonymous):

No ones "stupid" just others have more practice in that field than others and thanks @cwrw238

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

i got 6.22

OpenStudy (anonymous):

is that the correct answer??

OpenStudy (anonymous):

can u please help me i have until 8:00

OpenStudy (anonymous):

no subtract 3.00 from 6.25 so 6.25-3.00?

OpenStudy (anonymous):

i got 3.25

OpenStudy (anonymous):

yes now whats your new equation?

OpenStudy (anonymous):

should be in y=mx+b form